hi guys need your help on this one
A VENDOR SELLS HOT DOGS AND BAGS OF POTATO CHIPS. A CUSTOMER BUYS 2 DOGS AND 5 BAGS OF POTATO CHIPS FOR $12.00. ANOTHER CUSTOMER BUYS 5 HOT DOGS AND 3 BAGS OF POTATO CHIPS FOR $15.75. FIND THE COST OF EACH ITEM.

Answer by checkley77(12844)

2D+5C=12 MULTIPLY THIS EQUATION BY 5
5D+3C=15.75 MULTIPLY THIS EQUATION BY -2 & ADD.
10D+25C=60
-10D-6C=-31.5
----------------------------
19C=28.5
C=28.5/19
C=1.50 FOR EACH BAG OF CHIPE.
2D+5*1.50=12
2D+7.50=12
2D=12-7.50
2D=4.50
D=4.50/2
D=2.25 FOR EACH HOT DOG.
PROOF:
5*2.25+3*1.50=15.75
11.25+4.50=15.75
15.75